OpenStudy (anonymous):

If the sin 30° is 1 over 2, then the cos ____° = _____. 60° ; 1 over 2 30° ; square root 2 over 2 60° ; square root 3 over 2 30° ; 1

OpenStudy (anonymous):

I will give a medal please help

OpenStudy (anonymous):

sin is opp/hyp cos is adj/hyp

OpenStudy (whpalmer4):

Sin and Cos are complementary: Sin(x) = Cos(90-x) and Cos(x) = Sin(90-x)

OpenStudy (whpalmer4):

(where working in degrees, that is)

OpenStudy (anonymous):

would it be c

OpenStudy (anonymous):

opp = 1 adj = 2 hyp =

OpenStudy (anonymous):

but u said sin is opp over hyp so the opp would be 1 and the hyp would be 2

OpenStudy (whpalmer4):



OpenStudy (anonymous):

so it would be c right?

OpenStudy (anonymous):

yes I made a typo - you are right!

OpenStudy (whpalmer4):

 No. \[\sin(x) = \cos(90^\circ - x)\]

OpenStudy (whpalmer4):

If \(x = 30^\circ\), then \[\sin(30^\circ) = \cos(90^\circ-30^\circ) = \cos(60^\circ)\]

OpenStudy (anonymous):

but it would have to be 60 degrees and the other one with 60 degrees doesn't make sense

OpenStudy (whpalmer4):

The picture is symmetrical. The second one I drew shows the line for 60 degrees, and there the hypotenuse is 1 (because this is the unit circle), the adjacent side will be 1/2, and the opposite side will be sqrt(3)/2. Earlier, when the angle was 30 degrees, we had the hypotenuse = 1, the opposite side = 1/2 and the adjacent side = sqrt(3)/2.

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

60° ; 1 over 2 , because the angles are complementary i just took this test and got this wrong, the is the right answer